A Complete Bibliography of ACM Transactions on Evolutionary Learning and Optimization (TELO)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

30 April 2024 Version 1.03

Title word cross-reference

(1 + (λ, λ)) [FS22].

ACM [BW21]. Actions [CCKL22]. Activation [NA21]. Adaptation [MMM^{+}23, SB22]. Adaptive [YYY^{+}23]. Adversarial [BB23, HTAD^{+}21]. AI [BBC^{+}24]. Algorithm [FS22, JWT^{+}21]. Algorithms [BB23, CLOW21, DN21, DAB21, LBX21]. Allocation [CLOW21, MWD22]. Analyses [WVY^{+}22]. Analysis [AD21, DGNN22, FS22, MWD22, NRD^{+}24, Oli23, FCC23]. Application [AMM22, BW22, MMM^{+}23]. Approach [BZP^{+}24, DAB21]. Approximate [AMM22]. Approximation [MMM^{+}23]. Archiving [Li21]. Artificial [Li21]. Aspects [Mis24]. assisted [DAB21]. Attribution [WHLY24]. AutoML [PH22].

Based [CLOW21, RWF23, CS23, MSDG22, YYY^{+}23]. Bayesian [BW22, DERF21]. Behaviour [BZP^{+}24]. Berthing [AMM22, MMM^{+}23]. Best [CLOW21, FW23].

Capacitated [JWT^{+}21]. Cardinality [FKR^{+}23]. Case [MMM^{+}23]. Classifier [LBX21]. Code [RP22]. Coefficients [DAB21]. Coevolution [HTAD^{+}21, XLL^{+}21]. Combining [DAB21, RWF23, Sig23]. Compaction [LBX21]. Comparison [LBX21]. Computation [BBC^{+}24, KRT21, LIBP21].

Constrained [FKR^{+}23, XLL^{+}21]. Constraint [XLL^{+}21]. Constraint-Objective [XLL^{+}21]. Construction [JWT^{+}21]. Control [AMM22, FS22, MMM^{+}23]. Convergence [MSDG22]. Cooperative [XLL^{+}21]. Covariance [MMM^{+}23]. Creates [LRBW23]. Credible [BB23]. Crossover [FKR^{+}23]. Curiosity [LRBW23].

Damage [ASC^{+}23]. Data [LK21, RP22]. Deactivation [YYY^{+}23]. Decentralized [MWD22]. Deep [Lan22, Sig23]. Delay [LCI21]. Dependencies [DAB21]. Design [SB22]. Designing [DAB21]. Detailed [Oli23, WVY^{+}22]. Differential [TS23]. dimensional [BW22]. Discover [BZP^{+}24]. Discrete [DN21]. Diversity [ASC^{+}23, DGNN22, LRBW23, PMBR23]. Domains [FCC23]. Dominance [KRT21]. donor [WP22]. Dynamic [BZP^{+}24, MWD22, YYY^{+}23].

Editorial [BW21, FW23, GPC23]. Effectively [BZP^{+}24]. Effectiveness [PMBR23]. Efficient [DAB21, KRT21]. Electrical [BZP^{+}24]. Elites [FCC23]. Emergent [KSHB21, RP22]. Empirical [FS22, FCC23, NA21]. Energy [TS23]. Evolution [NA21, Sig23, SB22, TS23]. Evolutionary [BBC^{+}24, BZP^{+}24, BW21, CLOW21, DGNN22, DN21, GPC23, LIBP21, Mis24, MSDG22, MMM^{+}23]. Evolved [MWD22, PMBR23]. Evolving [RWF23]. Examples [BB23]. Explainability [BZP^{+}24]. Explainable [BBC^{+}24, FLP22, Mis24, WHLY24]. Explanation [WHLY24]. Exploitation [DERF21]. Exploration [DERF21]. Exploring [Mis24].

Factors [PMBR23]. Faster [MSDG22]. Feature [JWT^{+}21, WHLY24]. Forecasting [KSHB21]. Functions [AD21, LK21, NA21].

Gaussian [BW22]. GECCO [FW23]. General [SB22]. Generation [BB23]. Generative [HTAD^{+}21]. Genetic [BB23, DAB21, FS22, LK21, Lan22, LCI21, NRD^{+}24, RP22, WP22]. Good [DERF21]. Gradient [CS23]. Graphs [KSHB21]. Greed [DERF21].

heuristic [JWT^{+}21]. Heuristics

[WVY^{+}22]. Hierarchical [ASC^{+}23]. High [BW22]. High-dimensional [BW22]. Human [NRD^{+}24]. Human-in-the-loop [NRD^{+}24]. Hyperparameter [KPM^{+}23].

Impacting [PMBR23]. Improvement [LK21, LCI21, RP22]. Inaugural [BW21]. Ingredients [NRD^{+}24]. Innovized [MSDG22]. Input [Li21]. Interaction [Oli23]. Interpretable [NRD^{+}24]. Introduction [BBC^{+}24]. Inverse [CLOW21]. Investigation [NA21]. IOHanalyzer [WVY^{+}22]. Issue [BBC^{+}24, BW21, FW23, GPC23]. Iterative [WVY^{+}22].

Knowledge [LBX21].

Landscapes [PH22]. Large [XLL^{+}21]. Large-scale [XLL^{+}21]. Learnable [Mis24]. Learning [BW21, CCKL22, DAB21, GPC23, KPM^{+}23, LBX21, MSDG22, NRD^{+}24, RWF23, Sig23, TS23, WHLY24, WP22]. Learning-based [MSDG22]. Linear [BZP^{+}24]. loop [NRD^{+}24]. Loss [PH22].

Machine [KPM^{+}23, WHLY24]. Machines [BZP^{+}24]. Macro [CCKL22]. Manifolds [SB22]. MAP [FCC23]. Maths [LK21]. Matrix [MMM^{+}23, SB22]. Max [MMM^{+}23]. Mechanisms [FS22]. Meta [JWT^{+}21]. Meta-heuristic [JWT^{+}21]. Method [Mis24]. Methods [Li21]. Min [MMM^{+}23]. Min-Max [MMM^{+}23]. Minimization [AMM22]. Model [BZP^{+}24, CS23]. Model-based [CS23]. Modeling [BW22]. Models [NRD^{+}24]. Modular [PMBR23]. Multi [BZP^{+}24, KPM^{+}23, KRT21, MSDG22, WHLY24, WP22]. Multi-donor [WP22]. Multi-Objective [BZP^{+}24, KPM^{+}23, KRT21, MSDG22, WHLY24]. Multiobjective [Mis24, Li21]. Mutation [RWF23]. Mutation-Based [RWF23].

Navigation [KSHB21]. Network [HTAD^{+}21]. Networks [LCI21]. Neural [WP22]. Neuroevolution [FCC23]. Novel [DAB21].

Objective [BZP^{+}24, KPM^{+}23, XLL^{+}21, KRT21, MSDG22, WHLY24]. Observable [KSHB21]. offs [DERF21]. Online [ASC^{+}23, RWF23]. Operator [MSDG22]. Optimisation [DERF21]. Optimization [AMM22, BW22, DGNN22, DN21, FKR^{+}23, KPM^{+}23, KRT21, Mis24, MSDG22, MMM^{+}23, SB22, WVY^{+}22, XLL^{+}21, YYY^{+}23, BW21]. Oracle [AMM22]. Our [Li21]. Overview [KPM^{+}23].

P2P [TS23]. Parameter [FS22]. Part [FW23]. Partially [KSHB21]. Particle [YYY^{+}23]. Performance [Mis24, WVY^{+}22]. Permutation [CS23, DGNN22]. PGA [FCC23]. PGA-MAP-Elites [FCC23]. Physical [ASC^{+}23]. Plateau [AD21]. Point [AMM22]. Policy [LRBW23, Sig23]. Population [YYY^{+}23]. Precise [AD21]. Pressure [CLOW21]. Probabilistic [KRT21]. Problems [CS23, DGNN22, JWT^{+}21, YYY^{+}23]. Process [BW22]. Program [KSHB21]. Programming [Lan22, NRD^{+}24, WP22]. Progress [DN21, MSDG22]. Prospect [TS23]. Protocols [LCI21]. Prototypes [FLP22].

Quadratic [SB22]. Quality [ASC^{+}23]. Quality-Diversity [ASC^{+}23].

Rank [CLOW21]. Rank-Based [CLOW21]. Ranking [MMM^{+}23]. Rational [Oli23]. Recommendation [JWT^{+}21]. Recovery [ASC^{+}23]. Recursive [KSHB21]. Regression [BZP^{+}24, FLP22, NRD^{+}24, Oli23]. Reinforcement [CCKL22, GPC23, Sig23, TS23]. Reliable [Li21]. Representation [Oli23]. Reproducibility [LIBP21]. Reproductive [CLOW21]. Response [MWD22]. Results [Oli23]. Reusability [CCKL22]. Robots [ASC^{+}23, PMBR23]. Robust [AMM22, Lan22]. Routing [JWT^{+}21, LCI21]. Rule [LBX21]. Runtime [AD21].

Saddle [AMM22]. scale [XLL^{+}21]. Search [CS23, LRBW23, RWF23, Sig23]. Selective [CLOW21]. Sequences [Li21]. Simple [Li21]. Size [YYY^{+}23]. Software [RP22, RWF23]. Spatial [HTAD^{+}21]. Special [BBC^{+}24, FW23, GPC23]. Species [YYY^{+}23]. Species-based [YYY^{+}23]. SRBench [Oli23]. State [CLOW21]. Steady [CLOW21]. Steady-State [CLOW21]. Stochastic [RWF23]. Strategy [MMM^{+}23, SB22]. Surrogate [DAB21]. Surrogate-assisted [DAB21]. Survey [BW22, DN21, Sig23]. Swarm [YYY^{+}23]. Symbolic [NRD^{+}24, Oli23]. Synthesis [RP22]. System [LBX21]. Systems [RP22].

Tangled [KSHB21]. Task [MWD22]. Tasks [KSHB21, MMM^{+}23]. Theoretical [FS22]. Theory [DN21, TS23]. Thermal [BZP^{+}24]. Thresholds [MWD22]. Tolerant [LCI21]. Trade [DERF21]. Trade-offs [DERF21]. Tradeoffs [BZP^{+}24]. Trading [TS23]. Training [HTAD^{+}21]. Transactions [BW21]. Transfer [WP22]. Transferability [CCKL22]. Transformation [Oli23]. Transformation-Interaction-Rational [Oli23]. Trees [Lan22]. Trials [CLOW21].

Uncertain [FCC23]. Using [BZP^{+}24].

Vehicle [JWT^{+}21]. Via [FLP22]. Visualization [LBX21]. Visually [BB23]. ViZDoom [KSHB21].

Walsh [DAB21]. Worst [MMM^{+}23]. Worst-Case [MMM^{+}23].

References

Antipov:2021:PRA

[AD21] Denis Antipov and Benjamin Doerr. Precise runtime analysis for plateau functions. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1(4):13:1–13:28, December 2021. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3469800.

Akimoto:2022:SPO

[AMM22] Youhei Akimoto, Yoshiki Miyauchi, and Atsuo Maki. Saddle point optimization with approximate minimization oracle and its application to robust berthing control. ACM Transactions on Evolutionary Learning and Optimization (TELO), 2(1):2:1–2:32, March 2022. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3510425.

Allard:2023:ODR

[ASC^{+}23] Maxime Allard, Simón C. Smith, Konstantinos Chatzilygeroudis, Bryan Lim, and Antoine Cully. Online damage recovery for physical robots with hierarchical quality-diversity. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3 (2):6:1–6:??, June 2023. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic). URL

https://dl.acm.org/doi/10.
1145/3596912.

Bradley:2023:GVC

[BB23] James R. Bradley and A. Paul Blossom. The generation of visually credible adversarial examples with genetic algorithms. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3(1):2:1–2:??, March 2023. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3582276.

Bacardit:2024:ISI

[BBC^{+}24] Jaume Bacardit, Alexander Brownlee, Stefano Cagnoni, Giovanni Iacca, John McCall, and David Walker. Introduction to the special issue on explainable AI in evolutionary computation. ACM Transactions on Evolutionary Learning and Optimization (TELO), 4(1):1:1–1:??, March 2024. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3649144.

Branke:2021:ATE

[BW21] Juergen Branke and Darrell Whitley. ACM Transactions on Evolutionary Learning and Optimization inaugural issue editorial. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1(1):1e:1– 1e:2, June 2021. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic). URL

https://dl.acm.org/doi/10.
1145/3449277.

Binois:2022:SHD

[BW22] Mickaël Binois and Nathan Wycoff. A survey on highdimensional Gaussian process modeling with application to Bayesian optimization. ACM Transactions on Evolutionary Learning and Optimization (TELO), 2(2):8:1–8:??, June 2022. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3545611.

Banda:2024:MOE

[BZP^{+}24] Tiwonge Msulira Banda, Alexandru-Ciprian Zavoianu, Andrei Petrovski, Daniel Wöckinger, and Gerd Bramerdorfer. A multiobjective evolutionary approach to discover explainability tradeoffs when using linear regression to effectively model the dynamic thermal behaviour of electrical machines. ACM Transactions on Evolutionary Learning and Optimization (TELO), 4(1): 3:1–3:??, March 2024. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3597618.

Chang:2022:RTM

[CCKL22] Yi-Hsiang Chang, Kuan-Yu Chang, Henry Kuo, and Chun-Yi Lee. Reusability and transferability of macro actions for reinforcement learning. ACM Transactions on Evolutionary Learning and Optimization (TELO), 2(1): 4:1–4:16, March 2022. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3514260.

Corus:2021:SSE

[CLOW21] Dogan Corus, Andrei Lissovoi, Pietro S. Oliveto, and Carsten Witt. On steady-state evolutionary algorithms and selective pressure: Why inverse rankbased allocation of reproductive trials is best. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1(1):2:1–2:38, June 2021. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3427474.

Ceberio:2023:MBG

[CS23] Josu Ceberio and Valentino Santucci. Model-based gradient search for permutation problems. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3(4):15:1– 15:??, December 2023. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3628605.

Dushatskiy:2021:NAD

[DAB21] Arkadiy Dushatskiy, Tanja Alderliesten, and Peter A. N. Bosman. A novel approach to designing surrogate-assisted genetic algorithms by combining efficient learning of Walsh coefficients and dependencies. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1(2): 5:1–5:23, June 2021. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic). URL

https://dl.acm.org/doi/10.
1145/3453141.

DeAth:2021:GGE

[DERF21] George De Ath, Richard M. Everson, Alma A. M. Rahat, and Jonathan E. Fieldsend. Greed is good: Exploration and exploitation trade-offs in Bayesian optimisation. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1 (1):1:1–1:22, June 2021. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3425501.

Do:2022:AED

[DGNN22] Anh Do, Mingyu Guo, Aneta Neumann, and Frank Neumann. Analysis of evolutionary diversity optimization for permutation problems. ACM Transactions on Evolutionary Learning and Optimization (TELO), 2 (3):11:1–11:??, September 2022. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3561974.

Doerr:2021:SRP

[DN21] Benjamin Doerr and Frank Neumann. A survey on recent progress in the theory of evolutionary algorithms for discrete optimization. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1 (4):16:1–16:43, December 2021. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3472304.

Flageat:2023:EAP

[FCC23] Manon Flageat, Félix Chalumeau, and Antoine Cully. Empirical analysis of PGA-MAP-Elites for neuroevolution in uncertain domains. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3(1):1:1– 1:??, March 2023. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3577203.

Friedrich:2023:CCC

[FKR^{+}23] Tobias Friedrich, Timo Kötzing, Aishwarya Radhakrishnan, Leon Schiller, Martin Schirneck, Georg Tennigkeit, and Simon Wietheger. Crossover for cardinality constrained optimization. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3(2):5:1–5:??, June 2023. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3603629.

Filho:2022:ERP

[FLP22] Renato Miranda Filho, Anísio M.

Lacerda, and Gisele L. Pappa. Explainable regression via prototypes. ACM Transactions on Evolutionary Learning and Optimization (TELO), 2(4):14:1– 14:??, December 2022. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3576903.

Fajardo:2022:TEA

[FS22] Mario Alejandro Hevia Fajardo and Dirk Sudholt. Theoretical and empirical analysis of parameter control mechanisms in the (1 + (λ, λ)) genetic algorithm. ACM Transactions on Evolutionary Learning and Optimization (TELO), 2(4):13:1– 13:??, December 2022. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3564755.

Fieldsend:2023:EBG

[FW23] Jonathan Fieldsend and Markus Wagner. Editorial to the “Best of GECCO 2022” special issue: Part I. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3(2):4:1–4:??, June 2023. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3606034.

Gaier:2023:EER

[GPC23] Adam Gaier, Giuseppe Paolo, and Antoine Cully. Editorial to the “Evolutionary Reinforcement Learning” special issue. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3(3):9:1– 9:??, September 2023. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3624559.

Hemberg:2021:SCG

[HTAD^{+}21] Erik Hemberg, Jamal Toutouh, Abdullah Al-Dujaili, Tom Schmiedlechner, and Una-May O’Reilly. Spatial coevolution for generative adversarial network training. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1(2):6:1–6:28, June 2021. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3458845.

Jiang:2021:FCM

[JWT^{+}21] Hao Jiang, Yuhang Wang, Ye Tian, Xingyi Zhang, and Jianhua Xiao. Feature construction for meta-heuristic algorithm recommendation of capacitated vehicle routing problems. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1(1):3:1–3:28, June 2021. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3447540.

Karl:2023:MOH

[KPM^{+}23] Florian Karl, Tobias Pielok, Julia Moosbauer, Florian Pfisterer, Stefan Coors, Martin Binder, Lennart Schneider, Janek Thomas, Jakob Richter, Michel Lang, Eduardo C. Garrido-Merchán, Juergen Branke, and Bernd Bischl. Multi-objective hyperparameter optimization in machine learning — an overview. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3 (4):16:1–16:??, December 2023. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3610536.

Khosravi:2021:ECP

[KRT21] Faramarz Khosravi, Alexander Rass, and Jürgen Teich. Efficient computation of probabilistic dominance in multi-objective optimization. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1 (4):15:1–15:26, December 2021. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3469801.

Kelly:2021:ETP

[KSHB21] Stephen Kelly, Robert J. Smith, Malcolm I. Heywood, and Wolfgang Banzhaf. Emergent tangled program graphs in partially observable recursive forecasting and ViZDoom navigation tasks. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1(3):11:1– 11:41, September 2021. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3468857.

Langdon:2022:DGP

[Lan22] William B. Langdon. Deep genetic programming trees are robust. ACM Transactions on Evolutionary Learning and Optimization (TELO), 2(2):6:1– 6:??, June 2022. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3539738.

Liu:2021:CLC

[LBX21] Yi Liu, Will N. Browne, and Bing Xue. A comparison of learning classifier systems’ rule compaction algorithms for knowledge visualization. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1 (3):10:1–10:38, September 2021. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3468166.

Lorandi:2021:GIR

[LCI21] Michela Lorandi, Leonardo Lucio Custode, and Giovanni Iacca. Genetic improvement of routing protocols for delay tolerant networks. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1(1): 04:1–04:37, June 2021. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3453683.

Li:2021:OAR

[Li21] Miqing Li. Is our archiving reliable? Multiobjective archiving methods on “simple” artificial input sequences. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1(3):9:1–9:19, September 2021. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465335.

Lopez-ibanez:2021:REC

[LIBP21] Manuel López-Ibáñez, Juergen Branke, and Luís Paquete. Reproducibility in evolutionary computation. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1 (4):14:1–14:21, December 2021. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3466624.

Langdon:2021:GID

[LK21] William B. Langdon and Oliver Krauss. Genetic improvement of data for maths functions. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1(2):7:1–7:30, June 2021. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3461016.

LeTolguenec:2023:CCD

[LRBW23] Paul-Antoine Le Tolguenec, Emmanuel Rachelson, Yann Besse, and Dennis G. Wilson. Curiosity creates diversity in policy search. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3(3):12:1– 12:??, September 2023. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3605782.

Misitano:2024:EEA

[Mis24] Giovanni Misitano. Exploring the explainable aspects and performance of a learnable evolutionary multiobjective optimization method. ACM Transactions on Evolutionary Learning and Optimization (TELO), 4(1): 4:1–4:??, March 2024. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3626104.

Miyagi:2023:CMA

[MMM^{+}23] Atsuhiro Miyagi, Yoshiki Miyauchi, Atsuo Maki, Kazuto Fukuchi, Jun Sakuma, and Youhei Akimoto. Covariance matrix adaptation evolutionary strategy with worst-case ranking approximation for min-max optimization and its application to berthing control tasks. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3(2):

8:1–8:??, June 2023. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic). URL

https://dl.acm.org/doi/10.
1145/3603716.

Mittal:2022:LBI

[MSDG22] Sukrit Mittal, Dhish Kumar Saxena, Kalyanmoy Deb, and Erik D. Goodman. A learningbased innovized progress operator for faster convergence in evolutionary multi-objective optimization. ACM Transactions on Evolutionary Learning and Optimization (TELO), 2(1): 1:1–1:29, March 2022. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3474059.

Mathias:2022:AER

[MWD22] H. David Mathias, Annie S.

Wu, and Daniel Dang. Analysis of evolved response thresholds for decentralized dynamic task allocation. ACM Transactions on Evolutionary Learning and Optimization (TELO), 2(2):5:1–5:??, June 2022. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3530821.

Nader:2021:EAF

[NA21] Andrew Nader and Danielle Azar. Evolution of activation functions: an empirical investigation. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1(2):

8:1–8:36, June 2021. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic). URL

https://dl.acm.org/doi/10.
1145/3464384.

Nadizar:2024:AIL

[NRD^{+}24] Giorgia Nadizar, Luigi Rovito, Andrea De Lorenzo, Eric Medvet, and Marco Virgolin. An analysis of the ingredients for learning interpretable symbolic regression models with humanin-the-loop and genetic programming. ACM Transactions on Evolutionary Learning and Optimization (TELO), 4(1):5:1– 5:??, March 2024. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic). URL

https://dl.acm.org/doi/10.
1145/3643688.

OlivettiDeFranca:2023:TIR

[Oli23] Fabrício Olivetti De França. Transformation-interaction-rational representation for symbolic regression: a detailed analysis of SRBench results. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3 (2):7:1–7:??, June 2023. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic). URL

https://dl.acm.org/doi/10.
1145/3597312.

Pushak:2022:ALL

[PH22] Yasha Pushak and Holger Hoos. AutoML loss landscapes. ACM Transactions on Evolutionary Learning and Optimization (TELO), 2(3):10:1– 10:??, September 2022. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic). URL

https://dl.acm.org/doi/10.
1145/3558774.

Pigozzi:2023:FID

[PMBR23] Federico Pigozzi, Eric Medvet, Alberto Bartoli, and Marco Rochelli. Factors impacting diversity and effectiveness of evolved modular robots. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3(1):3:1–3:??, March 2023. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3587101.

Rainford:2022:CDS

[RP22] Penny Faulkner Rainford and Barry Porter. Code and data synthesis for genetic improvement in emergent software systems. ACM Transactions on Evolutionary Learning and Optimization (TELO), 2(2):7:1–7:??, June 2022. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3542823.

Renzullo:2023:ESC

[RWF23] Joseph Renzullo, Westley Weimer, and Stephanie Forrest. Evolving software: Combining online learning with mutation-based stochastic search. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3 (4):13:1–13:??, December 2023. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3597617.

Spettel:2022:DMA

[SB22] Patrick Spettel and Hans-Georg Beyer. On the design of a matrix adaptation evolution strategy for optimization on general quadratic manifolds. ACM Transactions on Evolutionary Learning and Optimization (TELO), 2(3):9:1– 9:??, September 2022. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3551394.

Sigaud:2023:CED

[Sig23] Olivier Sigaud. Combining evolution and deep reinforcement learning for policy search: a survey. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3(3):10:1– 10:??, September 2023. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3569096.

Timilsina:2023:PET

[TS23] Ashutosh Timilsina and Simone Silvestri. P2P energy trading through prospect theory, differential evolution, and reinforcement learning. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3 (3):11:1–11:??, September 2023. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3603148.

Wang:2024:MOF

[WHLY24] Ziming Wang, Changwu Huang, Yun Li, and Xin Yao. Multiobjective feature attribution explanation for explainable machine learning. ACM Transactions on Evolutionary Learning and Optimization (TELO), 4 (1):2:1–2:??, March 2024. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3617380.

Wild:2022:MDN

[WP22] Alexander Wild and Barry Porter. Multi-donor neural transfer learning for genetic programming. ACM Transactions on Evolutionary Learning and Optimization (TELO), 2(4):12:1– 12:??, December 2022. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563043.

Wang:2022:IDP

[WVY^{+}22] Hao Wang, Diederick Vermetten, Furong Ye, Carola Doerr, and Thomas Bäck. IOHanalyzer: Detailed performance analyses for iterative optimization heuristics. ACM Transactions on Evolutionary Learning and Optimization (TELO), 2(1):3:1–3:29, March 2022. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3510426.

Xu:2021:COC

[XLL^{+}21] Peilan Xu, Wenjian Luo, Xin Lin, Jiajia Zhang, Yingying Qiao, and Xuan Wang. Constraintobjective cooperative coevolution for large-scale constrained optimization. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1(3):12:1– 12:26, September 2021. CO-DEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3469036.

Yazdani:2023:SBP

[YYY^{+}23] Delaram Yazdani, Danial Yazdani, Donya Yazdani, Mohammad Nabi Omidvar, Amir H. Gandomi, and Xin Yao. A species-based particle swarm optimization with adaptive population size and deactivation of species for dynamic optimization problems. ACM Transactions on Evolutionary Learning and Optimization (TELO), 3(4):14:1–14:??, December 2023. CODEN ???? ISSN 2688-299X (print), 2688-3007 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3604812.